# Multi-Agent Q-Learning Dynamics in Random Networks: Convergence due to Exploration and Sparsity

### v1

@@ 3 Convergence of Q-Learning in Graphs
In this section, we analyse the relation between network topology and the convergence of (QLD). In particular, in Lemma 1 we determine a sufficient condition on the exploration rates $T_{k}$ such that (QLD) converges to the unique QRE with high probability as $N\rightarrow\infty$. This result has two components: (i) the equilibrium is unique, and (ii) it is asymptotically stable with respect to (QLD). In doing so, we show that the convergence of Q-Learning is strongly tied to the sparsity of edges in the underlying graph.
We parameterise the network structure by the parameters of the model from which they are drawn. In the case of the Erdős-Rényi model, this is the probability $p$ of each edge. For the Stochastic Block model, this will be the probabilities $p_{c}$ of edges within each community and the probability $q$ of edges between communities. To parameterise the payoffs, we refer to the intensity of identical interests introduced in Hussain et al. (2024).
@@
To prove our main result, we make the following assumption on the network polymatrix games.
@@
Each edge is assigned the same bimatrix game, i.e., $(A^{kl},A^{lk})=(A,B)$ for all $(k,l)\in\mathcal{E}$.
Assumption 1 is well motivated in the literature as it analyses the case where agents are engaged in the same interaction with multiple opponents, a format that is commonly used to study the emergence of cooperation (Zhang et al., 2014; Mukhopadhyay & Chakraborty, 2020) and congestion (Szabó & Szolnoki, 2015) in multi-agent systems. Indeed, the assumption of shared payoffs has led to a number of successes in studying large scale systems with many agents Perrin et al. (2020); Wu et al. (2024); Pérolat et al. (2022); Parise & Ozdaglar (2023); Hu et al. (2019). We adopt this approach, but also evaluate the validity of our claims outside of this assumption in our experiments (c.f. Figure 1).
@@
Let $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A})_{k},(A,B)_{(k,l)\in\mathcal{E}})$ be a network polymatrix game that satisfies Assumption 1. Also let $G$ be the adjacency matrix associated with the edge-set $\mathcal{E}$. If, for each agent $k$, the exploration rate $T_{k}$ satisfies
$T_{k}>\delta_{I}\rho(G),$ (3)
the QRE $\mathbf{x}^{*}$ of the game $\mathcal{G}$ is unique. In addition, for all initial conditions, trajectories of (QLD) converge to $\mathbf{x}^{*}$.
Notice that Lemma 1 makes two claims. The first is that the QRE is unique, whilst the second shows that it is asymptotically stable under (QLD). The combination of these statements ensures that there is a unique equilibrium solution that is learned by the agents.
@@ Discussion.
In fully connected networks, where all agents are interacting with one another, we have $\rho(G)=N-1$, which, by Lemma 1, implies that higher exploration rates are required to guarantee convergence as the number of agents increases. This agrees with previous findings in Sanders et al. (2018) and Hussain et al. (2023). However, recall from Sec. 2.2 that, as $T_{k}\rightarrow\infty$ for all agents $k$, the QRE tends towards the uniform distribution, a trivial solution in which payoffs are ignored and the game is played randomly.
From Lemma 1, it is clear that, by controlling each term in (3), (QLD) can converge even with reasonably low exploration rates [^1] The definition of ‘reasonable’ depends on the tolerable distance from the Nash Equilibrium on the part of the user. See (Hussain et al., 2024) for an upper bound on such a distance.. An example is the aforementioned pairwise zero-sum game in which $\delta_{I}=0$ and (QLD) is convergent for all $T_{k}>0$. This agrees with the result found by Leonardos et al. (2024). One may also control $\rho(G)$ by choosing networks in which the number of edges does not grow with the number of players. Such networks include lattices Szabó & Borsos (2016) or ring networks (Kleinberg et al., 2011), in which each agent $k$ is connected to agents $(k-1)\mod N$ and $(k+1)\mod N$ respectively. In this case, the agent has only two neighbours, regardless of $N$ and so the spectral radius is always $\rho(G)=2$.
In order to lift the restriction of a constant number of edges for each agent, we move towards the analysis of random networks, in which the probability of edge connections is parameterised. In doing so, we can make broader statements about the relationship between the convergence of (QLD) in many-player games and the underlying network. Our goal is to place an upper bound on the spectral radius term $\rho(G)$ in Lemma 1 in terms of the parameters of the random graph model.
@@
We begin with the Erdős-Rényi model (Erdös & Rényi, 2006), and show that the spectral radius is related to the probability of edge connections in the network. In the Erdős-Rényi model, a graph is generated by sampling each edge with a probability $p\in(0,1)$ independently. The adjacency matrix $G$ is then a random matrix with entries that are Bernoulli distributed with parameter $p$. Let the degree of a node $k$ be the number of edges $(k,l)\in\mathcal{E}$ that contain $k$. The expected degree of each node in a network drawn from the Erdős-Rényi model is therefore $p(N-1)$.
@@
Consider the symmetric adjacency matrix $G\in\mathbb{R}^{N\times N}$ of a network drawn from the Erdős-Rényi model. Then $G$ has a zero main-diagonal and the off-diagonal entries $\{[G]_{kl}\sim\texttt{Bernoulli}(p):1\leq k<l\leq N\}$ are independent, where $p>0$. Let $\sigma=\sqrt{p(1-p)}$. For any $\epsilon>0$, the following holds almost surely as $N\to\infty$
$\rho(G)\leq(N-1)p+(2\sigma+\epsilon)\sqrt{N}.$ (C1)
The bound is dominated by the term $p(N-1)$ as $N\rightarrow\infty$, which is exactly the expected degree of a node in the Erdős-Rényi model. We now connect this result with Lemma 1 to show the convergence of (QLD).
@@
Let $\epsilon>0$ and let $\mathcal{G}$ be a network polymatrix game satisfying Assumption 1 and suppose that the network is drawn from the Erdős-Rényi model. If
$T_{k}>\delta_{I}[(N-1)p+(2\sigma+\epsilon)\sqrt{N}],$
then it holds almost surely as $N\to\infty$ that all trajectories of (QLD) converge to a unique QRE $\mathbf{x}^{*}\in\Delta$.
Stochastic Block Model.
Networks are drawn from the Stochastic Block Model (Holland et al., 1983) by partitioning the nodes into $C$ communities. The probability of an edge between nodes $k$ and $l$ that are in the same community $c$ is given by $p_{c}\in(0,1)$, whereas the probability of an edge between nodes in different communities is given by $q\in(0,1)$. The adjacency matrix $G$ is then generated by sampling each edge independently according to Bernoulli distribution with probabilities $p_{c}$ and $q$. The expected degree of a node in community $c$ is $p_{c}(N_{c}-1)+q(N-N_{c})$, where $N_{c}$ is the number of nodes in community $c$. Notice that the Erdős-Rényi model can be seen as a limit case of the Stochastic Block Model for a single community, i.e., $N_{c}=N$. For simplicity, the following lemma assumes all communities are of equal size, specifically $N_{c}=N/C$ for all $c$. However, the proof stays valid even when the communites’ size varies.
Lemma 3.
Consider the symmetric adjacency matrix $G$ of a network drawn from the Stochastic Block Model. Then
$\begin{cases}[G]_{kl}\sim\texttt{Bernoulli}(p_{c}),&\text{ if }k\text{ and }l\in c,\\ [G]_{kl}\sim\texttt{Bernoulli}(q),&\text{ if }k\in c\text{ and }l\in c^{\prime}\neq c,\\ [G]_{kk}=0,&\text{ for }1\leq k\leq N,\end{cases}$
where the entries $\{[G]_{kl}:1\leq k<l\leq N\}$ are independent. For any $\epsilon>0$, it holds almost surely as $N\to\infty$
$\rho(G)\leq Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}+(2\sigma_{\max}+\epsilon)\sqrt{N},$ (4)
where $p_{\max}=\max\{p_{1},\ldots,p_{C}\}$ and $\sigma_{\max}=\max\{\sqrt{p_{1}(1-p_{1})},\ldots,\sqrt{p_{C}(1-p_{C})},\sqrt{q(1-q)}\}$.
@@
The generalizability of Lemmas 2 and 3 warrants clarification. First, the $\epsilon\sqrt{N}$ term can be replaced by $\epsilon N^{1/4+\delta}\ln N$ for any $\delta>0$, as can be seen in Appendix C.
By considering the $O(N)$ dominant term $Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}$ in the bound (4), we notice that it can be rearranged to $p_{\max}(\frac{N}{C}-1)+q(N-\frac{N}{C})$, which is exactly the expected degree of a node in a Stochastic Block Model with $p_{c}=p_{\max}$ and $N_{c}=N/C$ for all subcommunities $c$. We apply this in the following theorem to connect the probability of edge connections in the graph with the convergence of (QLD).
Theorem 2.
Let $\epsilon>0$ and let $\mathcal{G}$ be a network polymatrix game satisfying Assumption 1. Suppose that the network is drawn from the Stochastic Block model. Finally, let $p_{\max},\sigma_{\max}$ be defined as in Lemma 3. If
$T_{k}>\delta_{I}\left[Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}+(2\sigma_{\max}+\epsilon)\sqrt{N}\right],$
the trajectories of (QLD) converge to a unique QRE $\mathbf{x}^{*}\in\Delta$ almost surely as $N\to\infty$.
We conclude our results with a point of independent interest, in which we prove sufficient conditions on the uniqueness of the Nash Equilibrium in games played on networks drawn from the Erdős-Rényi and Stochastic Block models. These hold as corollaries of Theorems 1 and 2 and by recognising that the Nash Equilibrium and QRE coincide in the limit $T_{k}\rightarrow 0$ for all $k$.
Corollary 1.
Let $\epsilon>0$ and $\mathcal{G}$ be a network polymatrix game satisfying Assumption 1 and $G$ the adjacency matrix of the network. The Nash Equilibrium is unique almost surely as $N\to\infty$ if the network is drawn from the (i) Erdős-Rényi model and
$0>\delta_{I}\left[(N-1)p+(2\sigma+\epsilon)\sqrt{N}\right],$
where $\sigma,p$ are defined as in Lemma 2 or (ii) the Stochastic Block model and
$0>\delta_{I}\left[Nq+\frac{N}{C}(p_{\max}-q)-p_{\max}+(2\sigma_{\max}+\epsilon)\sqrt{N}\right],$
where $p_{\max},q,\sigma_{\max}$ are defined as in Lemma 3.
4 Convergence of Q-Learning in Random Graphs: experimental evaluation
Figure 1: Proportion of converged runs after simulating (QLD) for choices of $(p,T)$ in network games where the network is drawn from the Erdős-Rényi model. We depict the (top) Network Sato Game (middle) Network Shapley Game with $\beta=0.2$ (bottom) Conflict Network Game, each with varying $N$ across heatmaps. For the Network Sato Game ($\delta_{I}=0.2$), we also depict the lower bound predicted by Theorem 1 whilst for the Network Shapley Game $\delta_{I}=2.0$, the bound exceeds the range of the plot and, due to the random nature of the payoffs, it is not possible to depict the bound for the Conflict Network Game.
Figure 2: Variation of the empirical convergence boundary in the (Left) Network Sato Game (Right) Network Shapley Game as $N$ increases, for different values of $p$. The $y$-axis shows the smallest exploration rate for which all simulations converged. In the Network Sato Game, we also plot the theoretical bound as a dashed line.
Figure 3: Proportion of converged runs after simulating (QLD) for choices of $(p,T)$ in a Network Sato Game where the network is drawn from the Stochastic Block Model for varying choices of number of agents $N$ and $q$. The theoretical lower bound from Theorem 2 is plotted in white.
@@ 4 Convergence of Q-Learning in Random Graphs: experimental evaluation
The aim of this section is to understand the implications of Theorems 1 and 2 in games with a finite number of agents. To do this, we simulate the Q-Learning algorithm described in Sec. 2.3 on network games, where the network is drawn from the Erdős-Rényi or Stochastic Block models. We show that, as the sparsity of the network decreases, higher exploration rates are required to guarantee the convergence of Q-Learning, by finding a boundary between stable and unstable dynamics. We analyze how our lower bound and the true boundary evolve with $N$ and provide empirical evidence that our results may extend beyond the assumptions used to derive the theoretical bounds.
Remark.
In our experiments, we assume all agents have the same exploration rate $T$ and thus omit the $k$ notation. As our results are lower bounds across all $k$, this is without loss of generality.
@@
We first simulate Q-Learning in the Network Shapley and Network Sato Games, which are extensions of classical bimatrix games, introduced in Shapley (2016) and Sato et al. (2002) respectively, to the network polymatrix setting. In both cases, each edge defines the same bimatrix game, $(A^{kl},A^{lk})=(A,B)\;\forall(k,l)\in\mathcal{E}$ where in the Network Shapley Game:
@@
where $\epsilon_{X},\epsilon_{Y}\in\mathbb{R}$. In our experiments, we fix $\epsilon_{X}=0.5,\epsilon_{Y}=-0.3$ and $\beta=0.2$.
Previous works have shown that orbits of (QLD) in these games can exhibit complex behaviours such as limit cycles and chaos for low exploration rates.
@@
In Figure 1, we examine the convergence of Q-Learning in the case that the network is drawn from the Erdős-Rényi model. We vary the probability $p$ of edge connections and the exploration rate $T$ for different numbers $N$ of agents. For each parameter choice, we simulate Q-Learning for 3000 rounds and for 50 initial conditions and record whether the dynamics converged to a fixed point. We determine convergence by the variance of the agent strategies in the final 300 rounds of the simulation.
In the case of the Network Sato Game, we also depict the lower bound predicted by Theorem 1. We observe that the theoretical bound accurately models the increase in the exploration rate required for convergence as $p$ increases, except a small region as $p\rightarrow 1$. This is due to the $\sigma$ term in the bound, which vanishes as $N\rightarrow\infty$ and so the bound becomes dominated by the $p(N-1)$ term. We also observe that, as $N$ increases, the region of non-convergence expands rapidly for high $p$, whilst remaining relatively fixed for low $p$. This is consistent with previous results in the literature (Hussain et al., 2024; Parise & Ozdaglar, 2019). We expand on this in Figure 2 in which we plot the boundary as $N$ increases for different values of $p$. We observe that the rate at which the boundary grows increases with $p$.
Figure 1 also shows the influence of the $\delta_{I}$ term in the comparison of the Network Sato ($\delta_{I}=0.2$) and Network Shapley ($\delta_{I}=2.0$) Games. In particular, the boundary, for any given parameter choice, occurs at higher exploration rates in the Network Shapley Game.
Finally, we consider the Conflict Network Game (Ewerhart & Valkanova, 2020), where the payoff matrices need not be the same along each edge. A network polymatrix game is a conflict network game if each bimatrix game $(A^{kl},A^{lk})$ satisfies $(A^{kl})_{ij}=v^{k}(P^{kl})_{ij}-c^{kl}_{i}$ and $(A^{lk})_{ji}=v^{j}(P^{lk})_{ji}-c^{lk}_{j}$, where $v_{k},v_{l}>0$, $c_{kl}\in\mathbb{R}^{n_{k}},c_{lk}\in\mathbb{R}^{n_{l}}$ and $(P^{kl})_{ij}+(P^{lk})_{ji}=1$ for all $i\in S_{k},j\in S_{l}$. As pointed out in Ewerhart & Valkanova (2020), if $v_{k}=v_{l}$ for all $(k,l)\in\mathcal{E}$, the conflict network is equivalent to a network zero sum game. To generate Figure 1, we independently sample each $v_{k}$ from the uniform distribution over $[0,1]$. We then randomly generate The matrix $P^{kl}$ and subsequently generate $(A^{kl},A^{lk})$ such that the conflict network condition is satisfied. We find that, although the boundary is not as clearly defined due to the random nature of the payoffs, the implication that convergence occurs with lower exploration rates in sparser networks still holds.
@@
We repeat our experiments for the case that the network is drawn from a Stochastic Block Model. In Figure 3, we depict the proportion of converged experiments in the Network Sato Game, with Network Shapley depicted in the Appendix (Figure 4). Due to the large number of parameters in the Stochastic Block Model, we fix the number of agents in each community to $N_{c}=5$ and suppose that $p_{c}=p$ for all communities $c$. We then determine the convergence of (QLD) for different values of $p$ and $T$ and vary across plots $q$ and $N$. We observe that whilst each of these parameters plays a role in determining asymptotic behaviour, it is clear that convergence occurs less frequently in denser graphs, i.e., as $p$ and $q$ increase.

### v2

@@ 3 Convergence of Q-Learning in Graphs
In this section, we show that the convergence of (QLD) is closely related to the structure of the underlying graph. First, we present the problem setup: we specify the payoffs by the intensity of identical interests framework (Hussain et al., 2024) and specialize to a certain class of network polymatrix games, which satisfy Assumption 1 below. Second, we establish in Lemma 1 a sufficient condition on the exploration rates $T_{k}$ such that (QLD) converges to the unique QRE. We adapt this result to the random network setup using the Erdős-Rényi and Stochastic Block models. In both cases, we establish connections between (QLD) and the expected degree of a node in the network.
@@
We prove our main results under the following Assumption on network polymatrix games.
@@
Each edge is assigned the same bimatrix game, i.e., $(A^{kl},A^{lk})=(A,B)$ $\forall(k,l)\in\mathcal{E}$.
Note that Assumption 1 does not require $A^{kl}=A^{lk}$, rather it requires that each edge is associated with the same pair of matrices $(A,B)$. No assumption is made about which agent receives matrix $A$ and which receives matrix $B$. This is reflected in our experiments (Section 4) in which payoff matrices are randomly assigned to agents on each edge. This aspect of the setting is discussed in detail in Appendix D. In our experiments, we also validate that our claims outside of this Assumption by studying the Conflict Network Game (Ewerhart and Valkanova, 2020); see Section 4 for details.
@@
Let $\mathcal{G}=(\mathcal{N},\mathcal{E},(\mathcal{A})_{k},(A,B)_{(k,l)\in\mathcal{E}})$ be a network polymatrix game that satisfies Assumption 1, and let $G$ be the adjacency matrix associated with the edge-set $\mathcal{E}$. If for all agents $k\in\mathcal{N}$, $T_{k}>\delta_{I}\cdot\rho(G)$, then the QRE $\mathbf{x}^{*}$ of the game $\mathcal{G}$ is unique. Further, $\mathbf{x}^{*}$ is globally asymptotically stable under (QLD).
This result has two components: (i) the QRE is unique, and (ii) it is asymptotically stable with respect to (QLD). Together, these statements ensure that a unique equilibrium solution is learned.
@@ Discussion.
Recall from Section 2 that lower exploration rates are preferred, so that the equilibrium solution remains close to a Nash Equilibrium. Lemma 1 provides two quantities, $\delta_{I}$ and $\rho(G)$ which can be chosen appropriately to guarantee the convergence of (QLD). A common approach to address this is to fix the game to a specific class. As an example, restricting to pairwise zero-sum games, where $\delta_{I}=0$, ensures that (QLD) converges as long as $T_{k}>0$ for all $k\in\mathcal{N}$, regardless of $\rho(G)$. Our approach, in contrast, focuses on controlling the spectral radius $\rho(G)$, while treating $\delta_{I}$ to be fixed a-priori. To achieve this, we explore how network sparsity – a tunable parameter in decentralised systems such as robotic swarms or sensor networks – impacts the convergence of (QLD). Our analysis turns to random networks, where network sparsity is directly parameterised through the probability of edge connections. By deriving probabilistic upper bounds on the (random) spectral radius term $\rho(G)$ in terms of these parameters, we can directly link the network’s sparsity to the convergence or (QLD), providing robust convergence guarantees that hold for broad classes of networks.
We specify the network structure by the parameters of the model from which it is generated. In the Erdős-Rényi model, this is the probability $p$ that a pair of nodes is connected by an edge. In the Stochastic Block model, it is the set of within-community edge probabilities $p_{1},\ldots,p_{C}$ and between-community edge probability $q$. We stress that our theoretical bounds are not limited to these specific models: they apply to any random network model with independent entries (see Lemma 9).
@@
We begin with the Erdős-Rényi (ER) model (Erdös and Rényi, 2006), in which the graph is generated by independently sampling each edge with probability $p\in(0,1)$. The adjacency matrix $G$ is then a random matrix with entries that are Bernoulli distributed with parameter $p$. Let the degree of a node $k$ be the number of edges $(k,l)\in\mathcal{E}$ that contain $k$. The expected node degree in a network drawn from the ER model is therefore p(N1)p(N-1).
Stochastic Block Model.
Networks are drawn from the Stochastic Block (SB) model (Holland et al., 1983) by partitioning the nodes into $C$ communities. The probability of an edge between nodes $k$ and $l$ that are in the same community $c$ is given by $p_{c}\in(0,1)$, whereas the probability of an edge between nodes in different communities is $q\in(0,1)$. The adjacency matrix $G$ is then generated by sampling each edge independently according to the Bernoulli distribution with probabilities $p_{c}$ and $q$. The expected node degree in community $c$ is $p_{c}(N_{c}-1)+q(N-N_{c})$, where $N_{c}$ is the number of nodes in community $c$. Note that the higher $p_{c}$ and $q$, the higher the expected node degree, and further that the ER model is a particular case of the SB model for a single community, i.e., $N_{c}=N$.
We first place probabilistic bounds on the spectral radius of the adjacency matrices in the ER and SB models, which we then use in Theorems 1 and 2 to obtain the main convergence results. For simplicity, in the SB model case , the following results assume all communities are of equal size, specifically $N_{c}=N/C$ for all $c$. However, the proof stays valid when the communities’ sizes vary.
@@
Consider the symmetric adjacency matrix $G\in\mathbb{R}^{N\times N}$ of a network drawn from the ER model, i.e., $G$ has a zero main-diagonal and the off-diagonal entries $\{[G]_{kl}\sim\texttt{Bernoulli}(p):1\leq k<l\leq N\}$ are independent, where $p>0$. For any $\epsilon>0$, it holds with probability greater than $1-\epsilon$,
$\rho(G)\leq\rho^{ER}\vcentcolon=(N-1)p+\sqrt{2(N-1)\,p\,(1-p)\log{\frac{2N}{\epsilon}}}+\frac{2}{3}\log{\frac{2N}{\epsilon}}.$ (4)
Lemma 3.
Consider the symmetric adjacency matrix $G\in\mathbb{R}^{N\times N}$ of a network drawn from the SB model
[…]
where $\{[G]_{kl}:1\leq k<l\leq N\}$ are independent. For any $\epsilon>0$, it holds with probability greater than $1-\epsilon$
$\rho(G)\leq\rho^{SB}=\vcentcolon$ $(N-N/C)q+(N/C-1)p_{\text{max}}+$
$+$ $\sqrt{2\left((N-N/C)q(1-q)+(N/C-1)\sigma_{\text{p,max}}^{2}\right)\log{\frac{2N}{\epsilon}}}+\frac{2}{3}\log{\frac{2N}{\epsilon}}.$
where $p_{\max}=\max\{p_{1},\ldots,p_{C}\}$ and $\sigma_{p,\max}=\max\{\sqrt{p_{1}(1-p_{1})},\ldots,\sqrt{p_{C}(1-p_{C})}\}$.
Together with Lemma 1, the two results above allow us to establish a link between the edge connection probabilities in the graph – and thus the expected node degree – and the convergence of (QLD).
@@
Let $\epsilon>0$ and $\mathcal{G}$ a network polymatrix game satisfying Assumption 1 with corresponding network drawn from the ER model with parameter $p$. If
$T_{k}(N)>\rho^{ER}\text{ for all agents }k\in\mathcal{N},$
then with probability at least $1-\epsilon$, there is a unique QRE $\mathbf{x}^{*}\in\Delta$ and, for any initial condition, the (QLD) trajectories converge to $\mathbf{x}^{*}$.
Theorem 2.
Let $\epsilon>0$ and $\mathcal{G}$ a network polymatrix game satisfying Assumption 1 with corresponding network drawn from the SB model with probabilities $p_{1},\ldots,p_{C},q$. If
$T_{k}(N)>\rho^{SB}\text{ for all agents }k\in\mathcal{N},$
then with probability at least $1-\epsilon$, there is a unique QRE $\mathbf{x}^{*}\in\Delta$ and, for any initial condition, the (QLD) trajectories converge to $\mathbf{x}^{*}$.
@@
In both the ER and SB model, the dominant $O(N)$ terms in the spectral-radius bounds are governed by expected node degrees. Indeed, the leading term in $\rho^{ER}$, namely (N1)p(N-1)p, equals the expected degree in the ER model. Likewise, the leading term in $\rho^{SB}$, namely
$(N-N/C)q+(N/C-1)p_{\max},$
represents the maximum expected degree across all blocks in the SB network. Further, the sub-leading $O(\sqrt{N\log N})$ term is governed by the variances of the degree distributions: p(1p)p(1-p) in the ER case, and on the block-wise maximum variance $\sigma_{p,\max}$ and variance q(1q)q(1-q) in the SBM case.
Implications to learning in multi-agent games. Theorems 1 and 2 show that (QLD) converges with low exploration rates whenever the expected degree of nodes in the network is small. Importantly, previous works (Sanders et al., 2018) linked the onset of unstable dynamics to the total number of agents $N$, but only in the limited case where all players interact with one-another, i.e., $p=1$. Our bounds refine this finding by showing that the onset of instability is related to the expected degree in the network, not to $N$ explicitly. Further, once we control the expected degree, the variance terms are automatically controlled as well. As a consequence, the threshold for chaotic behavior increases from linear to exponential in $N$. In other words, with controlled expected degrees, chaotic dynamics can only emerge at exponentially large network sizes, far beyond the regimes considered in earlier analyses.
4 Experimental Evaluation
@@
The aim of this section is to illustrate the implications of Theorems 1 and 2. We simulate the Q-Learning algorithm on network games, where the network is drawn from the ER and SB models. Note that we simulate the discrete-time algorithm described in Section 2, rather than integrating the continuous-time model (QLD). This is to explore the agreement between the theoretical predictions and the discrete-time update, which is the algorithm that is applied in practice. Each experiment runs for 4000 iterations, and we assess convergence numerically, as explained in Appendix E.1. We show that, as the expected node degree of the network increases, higher exploration rates are required for Q-Learning to converge by finding the boundary between stable and unstable dynamics. We analyse how the boundary evolves with $N$ and provide empirical evidence that our results extend beyond the assumptions used to derive the bounds. In our experiments, we assume all agents have the same exploration rate $T$ and thus omit the dependency on $k$. As our results are lower bounds across all $k$, this is without loss of generality.
@@
We first simulate Q-Learning in the Network Shapley and Network Sato games, which are extensions of classical bimatrix games, introduced by Shapley (1964) and Sato et al. (2002) respectively, to the network polymatrix setting. In both cases, each edge defines the same bimatrix game, $(A^{kl},A^{lk})=(A,B)\;\forall(k,l)\in\mathcal{E}$ where in the Network Shapley game:
@@
where $\epsilon_{X},\epsilon_{Y}\in\mathbb{R}$. We fix $\epsilon_{X}=0.5,\epsilon_{Y}=-0.3$ and $\beta=0.2$.
We also study the Conflict Network game, proposed in (Ewerhart and Valkanova, 2020), which does not satisfy Assumption 1. In particular, each edge may be associated with a different bimatrix game. A network polymatrix game is a Conflict Network game if each bimatrix game $(A^{kl},A^{lk})$ satisfies $(A^{kl})_{ij}=v^{k}(P^{kl})_{ij}-c^{kl}_{i}$ and $(A^{lk})_{ji}=v^{j}(P^{lk})_{ji}-c^{lk}_{j}$, where $v_{k},v_{l}>0$, $c_{kl}\in\mathbb{R}^{n_{k}},c_{lk}\in\mathbb{R}^{n_{l}}$ and $(P^{kl})_{ij}+(P^{lk})_{ji}=1$ for all $i\in S_{k}$ and $j\in S_{l}$. To generate Conflict Network games, we independently sample each $v_{k}$ from the uniform distribution over $[0,1]$ and generate $P^{kl}$ by randomly sampling its elements from the uniform distribution over [5,5][-5,5]. We then construct $(A^{kl},A^{lk})$ such that the Conflict Network condition is satisfied.
Figure 1: Proportion of (QLD) simulations that diverge in network games with network drawn from the Erdős-Rényi model using varying exploration rates $T$, edge probabilities $p$, and numbers of agents $N$. Each heatmap uses $(T,p)$ values on the grid $0.05,4.25]\times[0.05,0.25]$. Convergence at low exploration rates is more likely in sparser networks, i.e., low pp. Further, the boundary between convergent and divergent behaviour shifts rapidly with the number of agents NN when $p$ is high, highlighting the need to control $p$ for large NN. Additional heatmaps are available in Appendix E.2.
Figure 2: Variation of the empirical convergence boundary in various network games as NN increases, for different values of $p$. The $y$-axis shows the smallest exploration rate for which all 50 simulations converged. Dense networks (high $p$) require substantially higher exploration rates to ensure convergence, while sparse networks (low $p$) maintain low convergence thresholds even for large NN. Recall that lower exploration rates are preferable so that the QRE is closer to the Nash equilibrium.
@@
In Figure 1, we examine the convergence properties of Q-Learning on networks generated from the Erdős-Rényi (ER) model. We vary the edge probability $p$, the exploration rate $T$, and the number of agents $N$. Figure 2 extends this analysis by plotting the convergence boundary as $N$ increases for different values of $p$. The key observation is that the rate of boundary growth increases with $p$, demonstrating that network density – parameterised by the edge probability $p$ – has a direct impact on learning dynamics. Our results reveal that dense networks (high expected node degree) require higher exploration rates to guarantee convergence, consistent with previous findings in (Sanders et al., 2018). Recall from Section 2, however, that higher exploration rates yield Quantal Response Equilibria (QRE) that deviate further from the Nash Equilibrium. Crucially, we find that this effect is substantially reduced in sparse networks (small $p$).
Figure 3: Proportion of (QLD) simulations that diverge in Network Sato games with network drawn from the Stochastic Block model, varying exploration rates $T$, intra-community edge probability $p$, inter-community edge probability $q$ and number of agents $N$. We use equidistant $(T,p)$ values in $0.05,4.25]\times[0.05,0.25]$. The convergence boundary is jointly controlled by $p$ and $q$, offering greater flexibility to achieve convergence in structured networks. Extra results are provided in Appendix E.2.
Figure 4: Probability density function of final strategy variation in Network Sato games on heterogeneous stochastic block networks with $N=200$ agents, showing the maximum strategy variation across agents during the final 300 iterations, computed from 1024 independent simulations. Networks contain five communities with varying intra-community connection probabilities $p$ (shown in legend) and fixed inter-community probability $q=0.1$. Communities with lower connectivity (blue, green) achieve convergence at lower exploration rates than densely connected communities (red, purple).
The Network Sato game in Figure 2 illustrates this phenomenon clearly. For $p=0.05$, (QLD) converges for $T<2.0$ even with over $150$ agents. Conversely, for $p=0.75$, convergence may fail with as few as 25 agents. These results also hold in the Conflict Network game, which extends beyond the setting of Assumption 1. Although there is some stochasticity in the boundary due to randomised payoffs, the fundamental trend remains: convergence occurs at lower exploration rates when the expected node degree is lower. These findings underscore the importance for practitioners of carefully designing network structures to ensure multi-agent learning produces stable, near-optimal equilibria.
@@
We next evaluate Q-Learning on networks generated from the Stochastic Block (SB) model. Figure 3 shows the proportion of diverged experiments in the Network Sato game, with results for the Network Shapley and Conflict games provided in Appendix E.2. Because the SB model has many parameters, we fix the number of agents in each community to $N_{c}=5$ and suppose that $p_{c}=p$ for all communities $c$. We then assess the convergence of (QLD) for different values of $p$ and $T$, while varying $q$ and $N$. We find that, although each parameter influences last-iterate behaviour, convergence occurs less frequently in denser graphs, i.e., as $p$ and $q$ increase. Finally, we examine the case in which $p_{c}$ varies across communities in Figure 4. We simulate $1024$ independent runs of QLD dynamics and compute the maximum difference between strategy components over the final iterations. The results show that communities with lower $p_{c}$ tend to exhibit smaller maximum differences. This suggests that allowing agents in different communities to use different exploration rates may yield convergence with smaller rates than required by Theorem 2.